40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1916     Entry time: Mon Aug 17 02:12:53 2009
Author: Yoichi 
Type: Summary 
Category: Computers 
Subject: FE bootfest 
Rana, Yoichi

All the FE computers went red this evening.
We power cycled all of them.
They are all green now.

Not related to this, the CRT display of op540m is not working since Friday night.
We are not sure if it is the failure of the display or the graphics card.
Rana started alarm handler on the LCD display as a temporary measure.
ELOG V3.1.3-